OpenStudy (anonymous):

change 6x+3y=12 in to point slope form

OpenStudy (anonymous):

You must rearrange the function, as it is currently in standard form. First move the 6x to the other side. The function becomes: 3y=12-6x. Now divide both sides by 3. You are left with: y=4-2x OR y=-2x+4. Hope that helped!

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

no problem :)